4gm of NAOH solid are dissolved in just enough water to make 1 litre of solution . What is the[H^+] of the solution
 

Grade:12

1 Answers

Arun
5 years ago
Mass of NaOH = 4g
Volume of solution = 1 L
To find: Concentration of Hydrogen ions
Solve: Molarity = Number of moles / Volume of solution
Number of moles == Mass / Molar mass
 Molar mass of NaOH = 40 g mol-1
Number of moles = 4/40
= 0.1 mole
Molarity = 0.1 / 1
= 0.1 M
[OH-] = 0.1 M
[H+] = Kw / [OH-]
[H+] = 10-14 / 0.1
[H+] = 10-13M